OpenStudy (anonymous):

Solve -4(x + 1) - 3 = -3(x - 4).

OpenStudy (anonymous):

-4x -4 -3 = -3x +12 solve for x

OpenStudy (anonymous):

hey ;) <3

OpenStudy (anonymous):

x - 4 - 3 = 12 ?

OpenStudy (anonymous):

x - 4 = 15 ? x = 19 ?

OpenStudy (anonymous):

-4x +3x = -x as 4 which is greater than 3 have -ve sign

OpenStudy (anonymous):

-4-3-12 = 4x -3x

OpenStudy (chosenmatt):

x=19

OpenStudy (anonymous):

Matt that is what I got and it was wrong.

OpenStudy (chosenmatt):

x=-19

OpenStudy (chosenmatt):

i mean

OpenStudy (anonymous):

Oh lol

OpenStudy (chosenmatt):

-19 not 19

OpenStudy (anonymous):

hey chosen matt

OpenStudy (anonymous):

How did you get -19? And not 19?

OpenStudy (anonymous):

so rude

OpenStudy (chosenmatt):

plug in -19 into the equation

OpenStudy (anonymous):

Awww I should have checked my answer...lol. Thanks for the help man!

OpenStudy (anonymous):

-4-3-12 = 4x -3x x = -19

OpenStudy (chosenmatt):

no problem and hey @RainbowBubbles01

OpenStudy (anonymous):

xD finally someone pays attention to me